OpenStudy (peterg2001):

help please In a group of students, the ratio of girls to boys is 3 to 2. If there are 15 girls, how many total students are there? 10 20 25 30

OpenStudy (anonymous):

Well, first, you divide 15 \(\div\) 3.

OpenStudy (anonymous):

What do you get?

OpenStudy (peterg2001):

@Directrix @ThisIsAdriAtHerBest

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (peterg2001):

@Directrix @ThisIsAdriAtHerBest

OpenStudy (anonymous):

Hai

OpenStudy (anonymous):

It's 10...

OpenStudy (anonymous):

3/2 =15/10

OpenStudy (peterg2001):

On a field trip, there are 12 adults and 14 students. What is the ratio of the number of adults to the total number of people on the field trip? 6 to 13 12 to 14 26 to 12 6 to 7

OpenStudy (anonymous):

12/14=6/7 dude just divide by 2

OpenStudy (peterg2001):

thanks but both were wrong

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

ohhh i see what i did wrong

OpenStudy (anonymous):

sorry o.o